# Resonance of an object floating within a surface wavefield

### v1

@@ IV CONCLUSION
We have investigated the interaction of a floating cylinder with surface waves in laboratory-scale experiments. In the first part, the bodies’ natural frequencies of oscillation were experimentally characterized and theoretically rationalized in the absence of external wave fields. We show that the resonance frequency of the heaving motion corresponds to the maximum of wave emission. The resonance frequency $f_{0}$ depends on the cylinder radius $R$ and its thickness $h$. We present a theoretical model that rationalizes the observed resonance frequencies for all the body dimensions characterized experimentally. We show that, in the high-frequency limit, the added mass of a floating body corresponds to half the added mass of the same body when it is fully submerged in the fluid bulk. Although this hypothesis is not fulfilled for heaving motions, the comparison with the experimental resonant frequency is quantitative. In the second part, we studied the effects of the floater vertical motion on the re-emitted wave field. We measure the wave field re-emitted by the floating cylinder and we show that, in the transverse direction, the re-emitted wave exhibits a minimum of emission for excitation frequency close to the body resonance frequency $f_{0}$. By comparing the wave fields re-emitted by a fixed and a free-floating cylinder with same dimension, we proved that the body oscillations are indeed responsible for the observed minimal transverse emission.
@@ I. The Disk Limit: An Uniform Approximation
Figure 8: Experimental data compared to our theoretical model, using the approximated formula (7). Each point in the plot represents a single experiment. Cylinder radii are depicted using a color code. For a given radius, cylinder heights are, from right to left in the plot : $h=5\,\rm{mm}$, $h=10\,\rm{mm}$, $h=15\,\rm{mm}$, $h=20\,\rm{mm}$.
One might then ask the question: for arbitrary $\rho,\rho_{w},R,h$, how well does $\Omega(\epsilon=0)$ do at approximating the natural frequency more generally when $\epsilon\neq 0$? To answer this question, we examine the ratio $\Omega(\epsilon=0)/\Omega(\epsilon)$. When this ratio is equal to unity, the approximation is exact. We compute.
@@
for all values of $\epsilon$. Remarkably, one may invoke the approximation outlined in (7) for any density ratio $r$, radius $R$, and height $h_{0}$, and be guaranteed that the error, as compared to the true solution of equation (5), is less than about 4%. Thus, equation (7) represents a very valuable approximation for estimating the natural vibration of a floating cylinder.
@@ II. Signal processing for profile extraction
The spatio-temporal profile reported in section III of the main text is processed in the Fourier space [39] to filter the waves reflected by the boundaries of the tank, and to mitigate spurious fluctuations associated with low wave numbers. This Fourier filtering process involved: transforming the profiles $\eta_{\parallel}(x,t)$ into the spatial and temporal frequencies domain using the Fourier transform $\mathcal{F}$; applying a filter function $H(k,\omega)$ to suppress undesired frequencies; transforming the filtered profiles back to the spatio-temporal domain using the inverse Fourier transform $\mathcal{F}^{-1}$. These steps can be written as:
@@
Where, $\hat{\eta}_{\parallel}(k,\omega)$ corresponds to the Fourier transformation of the horizontal profile and, $\tilde{\eta}_{\parallel}(x,t)$ represents the filtered spatio-temporal signal. This filtered signal is then demodulated at the excitation frequency $f$, leading to a function $\tilde{\eta}_{\parallel}^{f}(x,t)$. A frame of this demodulated profile is represented as the blue curve in figure 5e in the main text. In order to measure the incoming wave amplitude, we extract the envelope of this demodulated profile via a peak detection algorithm that detects the local maxima of the demodulated curve $\tilde{\eta}_{\parallel}^{f}(x,t)$ for all time steps. The envelope curve $E_{\parallel}(\tilde{\eta}_{\parallel}^{f})$ is represented in red in figure 5e in the main text.
@@
First, a filter $H_{r}(k,\omega)$ is applied to the transverse profile in the Fourier space to filter the waves reflected by the boundaries of the tank. Then, Butter-worth filter function $H_{bw}(k,\omega)$ is applied to suppress undesired frequencies at low and high wave numbers [40, 41]. This filtered signal is then transformed back to a spatiotemporal signal $\tilde{\eta}_{\perp}(y,t)$ using inverse Fourier transform $\mathcal{F}^{-1}$ :
@@
Finally, the profile $\tilde{\eta}_{\perp}(y,t)$ is demodulated at the excitation frequency $f$ to get the spatio-temporal signal $\tilde{\eta}_{\perp}^{f}(y,t)$, plotted as the blue curve in figure 5f in the main text. The maximum amplitude of the waves emitted by the object in this direction, is obtained by extracting the envelope of $\tilde{\eta}_{\perp}^{f}(y,t)$, as described in the previous section. The envelope $E_{\perp}(\tilde{\eta}_{\perp}^{f})$ is represented in red in figure 5f in the main text.
III. Propagation waves of the cylinders for a fixed height
Using the same approach as in Figure 6 a of the main text, the ratio $A_{o}/A_{w}$ was determined for different heights of the free-floating cylinders with fixed radius $R=15\>\rm mm$. The results are plotted in Figure 9 as a function of the excitation frequency $f$ normalized by the measured resonance frequency $f_{0}$ of each body. Here also the system exhibits a minimum value of the ratio $A_{o}/A_{w}$ at an excitation frequency close to the resonance frequencie. Therefore, near its resonance frequency $f_{0}$, a cylindrical object shows a reduced emission efficiency in the direction orthogonal to the initial wave propagation.

### v2

@@ IV CONCLUSION
We have investigated the interaction of a floating cylinder with surface waves in laboratory-scale experiments. In the first part, the bodies’ natural oscillation frequencies were experimentally characterized and theoretically rationalized without external wave fields. We show that the resonance frequency of the heaving motion corresponds to the maximum wave emission. The resonance frequency $f_{0}$ depends on the cylinder radius $R$ and its thickness $h$. We present a theoretical model that rationalizes the observed resonance frequencies for all the body dimensions characterized experimentally. We show that, in the high-frequency limit, the added mass of a floating body corresponds to half the same body’s added mass when fully submerged in the fluid bulk. Although this hypothesis is not fulfilled for heaving motions, the comparison with the experimental resonant frequency is quantitative. In the second part, we studied the effects of the floater’s vertical motion on the re-emitted wave field. We measured the wave field re-emitted by the floating cylinder and showed that, in the transverse direction, the re-emitted wave exhibits a minimum emission for an excitation frequency close to the body resonance frequency $f_{0}$. By comparing the wave fields re-emitted by a fixed and a free-floating cylinder with the same dimension, we proved that the body oscillations are indeed responsible for the observed minimal transverse emission.
@@ I. The Disk Limit: An Uniform Approximation
Figure 8: Experimental data compared to our theoretical model, using the approximated formula (7). Each point in the plot represents a single experiment.Cylinder radii follow the color bar on the right. For a given radius and from right to left in the plot, cylinder heights are: $h=5\,\rm{mm}$, $h=10\,\rm{mm}$, $h=15\,\rm{mm}$, $h=20\,\rm{mm}$.
One might then ask the following question: for arbitrary $\rho,\rho_{w},R,h$, how well does $\Omega(\epsilon=0)$ approximates the natural frequency more generally when $\epsilon\neq 0$? To answer this question, we examine the ratio $\Omega(\epsilon=0)/\Omega(\epsilon)$. When this ratio is equal to unity, the approximation is exact. We compute
@@
for all values of $\epsilon$. Remarkably, one may invoke the approximation outlined in (7) for any density ratio $r$, radius $R$, and height $h_{0}$ and be guaranteed that the error, as compared to the solution of equation (5), is less than about 4%. Thus, equation (7) represents a very valuable approximation for estimating the natural vibration of a floating cylinder.
@@ II. Signal processing for profile extraction
The spatiotemporal profile reported in section III of the main text is processed in the Fourier space [39] to filter the waves reflected by the tank’s boundaries and mitigate spurious fluctuations associated with low wave numbers. This Fourier filtering process involved transforming the profiles $\eta_{\parallel}(x,t)$ into the spatial and temporal frequencies domain using the Fourier transform $\mathcal{F}$; applying a filter function $H(k,\omega)$ to suppress undesired frequencies; transforming the filtered profiles back to the spatio-temporal domain using the inverse Fourier transform $\mathcal{F}^{-1}$. These steps can be written as:
@@
where $\hat{\eta}_{\parallel}(k,\omega)$ corresponds to the Fourier transform of the horizontal profile and $\tilde{\eta}_{\parallel}(x,t)$ represents the filtered spatio-temporal signal. This filtered signal is then demodulated at the excitation frequency $f$, leading to the function $\tilde{\eta}_{\parallel}^{f}(x,t)$. A frame of this demodulated profile is represented as the blue curve in figure 5e in the main text. To measure the incoming wave amplitude, we extract the envelope of this demodulated profile via a peak detection algorithm that detects the local maxima of the demodulated curve $\tilde{\eta}_{\parallel}^{f}(x,t)$ for all time steps. The envelope curve $E_{\parallel}(\tilde{\eta}_{\parallel}^{f})$ is represented in red in figure 5e in the main text.
@@
First, a filter $H_{r}(k,\omega)$ is applied to the transverse profile in the Fourier space to filter the waves reflected by the boundaries of the tank. Then, a Butter-worth filter function $H_{bw}(k,\omega)$ is applied to suppress undesired frequencies at low and high wave numbers [40, 41]. This filtered signal is then transformed back to a spatiotemporal signal $\tilde{\eta}_{\perp}(y,t)$ using the inverse Fourier transform $\mathcal{F}^{-1}$:
@@
Finally, the profile $\tilde{\eta}_{\perp}(y,t)$ is demodulated at the excitation frequency $f$ to obtain the spatio-temporal signal $\tilde{\eta}_{\perp}^{f}(y,t)$, plotted as the blue curve in figure 5f in the main text. The maximum amplitude of the waves emitted by the object in this direction is obtained by extracting the envelope of $\tilde{\eta}_{\perp}^{f}(y,t)$, as described in the previous section. The envelope $E_{\perp}(\tilde{\eta}_{\perp}^{f})$ is represented in red in figure 5f in the main text.
III. Effect on wave field from cylinders with fixed height
Using the same approach as in Figure 6 of the main text, the ratio $A_{o}/A_{w}$ was determined for different heights of the free-floating cylinders with fixed radius $R=15\>\rm mm$. The results are plotted in Figure 9 as a function of the excitation frequency $f$ normalized by the measured resonance frequency $f_{0}$ of each body. Here, the system also exhibits a minimum value of the ratio $A_{o}/A_{w}$ at an excitation frequency close to the resonance frequency. Therefore, near its resonance frequency $f_{0}$, a cylindrical object shows a reduced emission efficiency in the direction orthogonal to the initial wave propagation.